State true or false the function f defined by f (x)=| x − 2| is differentiable in [0 ,1 ].
1
Expert's answer
2020-10-16T14:07:32-0400

The function is differentiable in [0 ,1] when it has derivative in every point in [0 ,1].

"f'(x)=(|x-2|)'" can be rewritten as

"f'(x)=(\sqrt{(x-2)^2})'=\frac{((x-2)^2)'}{2\sqrt{(x-2)^2}}="

"=\frac{2(x-2)}{2\sqrt{(x-2)^2}}=\frac{x-2}{|x-2|}" , where "x-2\neq0". So the function "f(x)=|x-2|" hasthe derivative in every point except "x=2" .Answer: true, the function "f(x)=|x-2|" is differentiable in [0 ,1].